MICQ(7) MICQ(7) mICQ− ,mICQ.,,,.:([]),(<>)().,<uin>UIN,<>,,UIN.,. , msg<>[<>] .,(\),.,(.).(#). −.{{{}}},.{%{}%},(GnuTLSOpenSSL).{<<>>}icq8(−2). ,<<<>>>icq5,−1,−4.. ;,.,,,−−2.,.,,,,−4. ,UTF−8.,,,,.−1,−4,,,. a[<>],..msg. r[<>]()..msg.

url<><><>

,. sms[<>][<>]<> SMS,+</><>,,.,.-. getauto[auto|away|na|dnd|occ|ffc][<>] ,away,nadndoccffc.auto,..0.4.10. auth[req|grant|deny|add]<> ,. resend<> . last[<>] ,,,

h<>[<N>[<>]]

history<>[<N>[<>]] <>.N−-N,N,20-,−,20. historyd<><>[<>] <>.,ISO8601,−−,−−::,,20. find<><> ,. finds<><> . tabs,. , loginconnlogin. online"". away[<>]

""..

na[<>]

""..

occ[<>]

""..

mICQ 1 MICQ(7) MICQ(7) dnd[<>]

""..

ffc[<message>]

""..

inv"". change[<>[<>]] ...

f<>

finger<>

<>.
ss<><>.

i. s[<>],,,. e,ee,eg,eeg,ev,eev,egv,eegv,w,ww,wg,wwg,wv,wwv,wgv,wwgv .: +,.ww. #,UIN’-.www. *,. −. ˆ :,. ww*ee*.,,ˆ . &. |. :. ˆ ,IP. ww*ee*,,: S. s,. .,."". ´´ ,. e[<>],.. ee[<>],,. eg,eeg ,.0.4.10. w,ww,wg,wwg ,. ev,eev,egv,eegv,wv,wwv,wgv,wwgv ,(shadow).0.5. ewide,,. wide. mICQ 2 MICQ(7) MICQ(7) search search<@> search<> search<><> @,(@),.,,,,. rand[<>] .,. add<uin><> uin. :,save.

add<><>

addalias<><> .addalias0.4.10. :,,save. add[<>[<>] addgroup<>[<>] ,,,addgroup.0.4.10. :,,save. rem[all]<> remalias[all]<> .,all,. (remalias)0.4.10. :,,save. remall<>

rem<><>

remgroupall<> remgroup<><> ,,all.0.4.10. :,,save. togig<> /. toginv<> /,. togvis<> /,. ICQ pass<> ICQ. :o´ (0xf3). :save,˜ /.micq/micqrc(.micqrc(7)),. update (e-mail,,). other,,. about(,). mICQ 3 MICQ(7) MICQ(7) setr[<>] ... reg<>. MICQ verbose[<>] mICQ.-: 8. 16. 32. 64. 256v6. 4096v7/v8. 8192v7/v8. 16284. 65536. 131072 . 262144 . 2097152 . 4194304 /. ,1,. clear. sound[on|off|] −on,−off,.0.4.10. autoaway[<>|on|off] ,,on−,off−.0off.,. alias[<><>] ,.,.%s,,.0.4.10. unalias<> .0.4.10. lang[<>|<N>]... ,ISON.,LANG,LCALLLCMESSAGES,!,autodefault.,.,none unload.,debug. ˜ /.micq/i18n/*.i18n,BASE/i18n/*.i18b˜ /.micq/i18n/*.i18n,,"",funny. ,,,117,"transde117default". uptime ,mICQ,-.

set<><>

on()off,().: mICQ 4 MICQ(7) MICQ(7) color:, delbs:delete()backspace(). funny: , auto:,, uinprompt: , autosave: micqrc, autofinger: UIN, linebreak: simple()−,,break()−,indent()−,smart()−,; 0.4.10.hermit,log,logonoff,silent,tabs0.5;opt. opt[<>|<>||global[<>[<>]]] ,,,,. ,,,,−,,−,,−.,0,−FALSE(),"nocolor"(),−. ,,,−. ,: colorscheme .,,0. colornone ,.black,red,green,yellow,blue,magenta,cyan,white,none,bold ANSI.,none,bold.,. colorserver ,. colorclient ,. colorinvchar ,,. colorerror ,. colordebug . colorquote . webaware . hideip IP. dcauth IP. dccont IP. mICQ 5 MICQ(7) MICQ(7) obeysbl/ .,. ignore/ . hidefrom/ . intimate/ . logonoff/ . logchange/ . logmess/ . showonoff/ . showchange/ . hideack/ . :,,. wantsbl/ . shadow/ ,*v. local/ . encoding/ ,,,,-. tabspoolinteger . autoaway/

"",.

autona/

"",.

autoocc/

"",.

autodnd/

"",.

autoffc/

"",.

colormessage/ . colorsent/ . mICQ 6 MICQ(7) MICQ(7) colorack/ . colorincoming/ . colorcontact/ . 0.5. optcontact<>[<>[<>]] .0.5. optgroup<>[<>[<>]] .0.5. optconnection[<>[<>]] .0.5. optglobal[<>[<>]] .0.5. save˜ /.micq/micqrc,.,. q quit exitmICQ.quitexit0.4.10. xmICQ.0.5. meta[show|load|save|set|get|rget]<> -.-: show-. load-. save-. set-. get-. getr-,,. 0.4.10. file[...] peerfile.0.4.10. accept[...] peeraccept.0.4.10.4. peer<><uin|nick> ,uinnick. openTCP. close. off,.

file<><>

. mICQ 7 MICQ(7) MICQ(7) files[<><as>]...<> .,.as´ /´ ,,´ .´ ,. accept[<>][<id>] ,id.,id,,. deny[<>][<id>][<>] id,. login[...] connlogin.0.4.10.4. conn[<><>] ,. open. login. close.. remove (). select.UIN,. contact[<>] : show. diff(/UIN),. add. upload . :,wantsbl.,optglobalwantsblon. peek<> .IRC.. :.

as<|uin><>

mICQ,uin. tclscript<> tcl,mICQ. tcl<>tcl.tclmicqhelptcl,mICQ. micq(1),micqrc(5) <ja2morrison@student.math.uwaterloo.ca>mICQ... mICQ 8